# Performance and radiation damage mitigation strategy for silicon photomultipliers on LEO space missions

### v4

@@ 6 Radiation damage in silicon with proton irradiation
After the full characterisation of the sensors in section 2, we have irradiated them with protons and electrons in order to estimate the radiation damage. When an energetic charged particle hits a SiPM, it deposits energy through both ionizing and non-ionizing processes. Radiation damage of SiPMs has been extensively discussed in [40]. In silicon, bulk damage is due to Non-Ionizing Energy Loss (NIEL) and surface damage by Ionizing Energy Loss (IEL). Bulk damage is due to high-energy particles (can be protons, neutrons, electrons, pions or photons), which can displace atoms out of their lattice creating defects as shown in figure 14. At energies beyond the keV scale, cluster defects can form when more atoms are displaced. Electrons beyond energies of 1 MeV typically can produce single defects, whereas heavier particles like protons also produce cluster defects. In the NIEL hypothesis, the radiation damage is proportional to the non-ionizing energy loss of the particles and this energy loss is proportional to the energy used to dislocate the lattice atoms (displacement energy). The defects in silicon crystals lead to an increase in leakage current due to the generation of electron-hole pairs from defects in the depletion region and trap-assisted tunnelling. This leads to a temperature-dependent increase in DCR, which is more significant for protons than for electrons, which primarily induce ionisation rather than defects due to their lower mass and less efficient energy transfer to the crystal lattice, resulting in a minimal increase in DCR. A detailed computation of the NIEL for protons and electrons in silicon can be done with the Screened Relativistic (SR) Treatment for NIEL Dose Nuclear and Electronic Stopping Power Calculator [41]. Hence, we expect that the damage caused by electrons induce a lower increase in DCR when compared to protons at equal over-voltage, temperature, cell sizes, and dose.
@@ 9 DCR prediction during mission
The corresponds to the measured current divided by , namely scaling for the ratio of areas to take into account the different SiPM sizes. The exponential term provides the measured current at desired temperature following the model in equation 7.1. Combining the DCR results with the dose-per-time response from the simulation yields the results in figure 22. As can be seen from the graph, the damage caused by electrons is very low when compared to protons at equal over-voltage, temperature, cell sizes, and dose, as expected. Specifically, 2 Gy of electrons corresponds to a DCR of approximately 0.15 MHz, while 2 Gy from protons results in a DCR of around 50 MHz, starting from the same DCR at 0 Gy.
@@ 9.1 Annealing
In the laboratory, we study the annealing effect by placing the irradiated devices in a climate chamber at a fixed temperature for a certain period and regularly measuring their IV curve. Then we measure again the current and compare it with the SiPM current before irradiation at the given over-voltage of 10 V to determine the recovery efficiency of the annealing process. The time evolution of the 10 V over-voltage current for different storage temperatures is shown in figure 23 for three types of SiPMs. In this figure, the y-axis shows the normalised current, meaning the dark current relative to that measured during the first measurement after irradiation. It can be noticed that the current is following an exponential decay with time, whose amplitude, time constant, and offset varying with the storage temperature. At time h, the measurement immediately after irradiation is shown. At time in the plot corresponds the first annealing measurement. This takes into account the effect of annealing without temperature to which the sensors were subjected during transport from the irradiation site. The data points in figure 23 are therefore fitted with the following function:
@@
Using this information, we constructed a model to estimate the expected DCR in space due to accumlating irradiation and the planned annealing processes. These two components change the DCR during the mission and can be described by the following differential equation:
@@ 10 Conclusions
This work presents the characterisation of FBK SiPMs, which allowed to select the optimal size of the -cells of SiPMs to adopt in the Terzina mission, and the study that determines the power consumption increase with radiation damage. It also lead to the definition of a mitigation strategy based on periodic annealing during the mission. A model has been developed from the measurements which allows other users to determine similar mitigation strategies to this relevant damaging effect of SiPMs.

### v5

@@ 6 Radiation damage in silicon with proton irradiation
After the full characterisation of the sensors in section 2, we have irradiated them with protons and electrons in order to estimate the radiation damage. When an energetic charged particle hits a SiPM, it deposits energy through both ionizing and non-ionizing processes. Radiation damage of SiPMs has been extensively discussed in [39]. In silicon, bulk damage is due to Non-Ionizing Energy Loss (NIEL) and surface damage by Ionizing Energy Loss (IEL). Bulk damage is due to high-energy particles (can be protons, neutrons, electrons, pions or photons), which can displace atoms out of their lattice creating defects as shown in figure 14. At energies beyond the keV scale, cluster defects can form when more atoms are displaced. Electrons beyond energies of 1 MeV typically can produce single defects, whereas heavier particles like protons also produce cluster defects. In the NIEL hypothesis, the radiation damage is proportional to the non-ionizing energy loss of the particles and this energy loss is proportional to the energy used to dislocate the lattice atoms (displacement energy). The defects in silicon crystals lead to an increase in leakage current due to the generation of electron-hole pairs from defects in the depletion region and trap-assisted tunnelling. This leads to a temperature-dependent increase in DCR, which is more significant for protons than for electrons, which primarily induce ionisation rather than defects due to their lower mass and less efficient energy transfer to the crystal lattice, resulting in a minimal increase in DCR. A detailed computation of the NIEL for protons and electrons in silicon can be done with the Screened Relativistic (SR) Treatment for NIEL Dose Nuclear and Electronic Stopping Power Calculator [40]. Hence, we expect that the damage caused by electrons induces a lower increase in DCR when compared to protons at equal over-voltage, temperature, cell sizes, and dose.
@@ 9 DCR prediction during mission
The corresponds to the measured current divided by , namely scaling for the ratio of areas to take into account the different SiPM sizes. The exponential term provides the measured current at the desired temperature following the model in equation 7.1. Combining the DCR results with the dose-per-time response from the simulation yields the results in figure 22. As can be seen from the graph, the damage caused by electrons is very low when compared to protons at equal over-voltage, temperature, cell sizes, and dose, as expected. Specifically, 2 Gy of electrons corresponds to a DCR of approximately 0.15 MHz, while 2 Gy from protons results in a DCR of around 50 MHz, starting from the same DCR at 0 Gy.
@@ 9.1 Annealing
In the laboratory, we studied the annealing effect by placing the irradiated devices in a climate chamber at a fixed temperature for a certain period and regularly measuring their IV curve. Then we measured again the current and compare it with the SiPM current before the annealing at the given over-voltage of 10 V to determine the recovery efficiency of the annealing process. The time evolution of the 10 V over-voltage current for different storage temperatures is shown in figure 23 for three types of SiPMs. In this figure, the y-axis shows the normalised current, meaning the dark current relative to that measured during the first measurement after irradiation. It can be noticed that the current is following an exponential decay with time, whose amplitude, time constant, and offset vary with the storage temperature. At time h, the measurement taken immediately after irradiation is shown (normalised current is equal to 1). In the plot, the first annealing measurement corresponds to time . This takes into account the effect of the sensors being subjected to annealing without heating during transport from the irradiation site. The data points in figure 23 are therefore fitted with the following function:
@@
Using this information, we constructed a model to estimate the expected DCR in space due to accumulating irradiation and the planned annealing processes. These two components change the DCR during the mission and can be described by the following differential equation:
@@ 10 Conclusions
This work presents the characterisation of FBK SiPMs, which allowed to select the optimal size of the -cells of SiPMs to adopt in the Terzina mission, and the study that determines the power consumption increase with radiation damage. It also leads to the definition of a mitigation strategy based on periodic annealing during the mission. A model has been developed from the measurements which allows other users to determine similar mitigation strategies to this relevant damaging effect of SiPMs.